Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 1974-01-24
Repair and maintenance accomplishments of the three manned Skylab missions.

Significant repair and maintenance accomplishments of the three manned Skylab missions were reported.

For the first 11 days, ground operations control kept the Skylab alive and allowed time for planning and hardware manufacturing necessary to salvage the mission. The first manned crew deployed the parasol which brought the OWS internal temperatures down. On the 13th mission day, beam 1 and its solar array were deployed to provide adequate power for normal mission. The crew succeeded in reactivating a malfunctioning battery, providing additional power. (This was accomplished by whacking it with a hammer during EVA.)

The first crew performed a variety of repairs and fixes on experiments, such as disassembly and repair of the S019 (UV stellar astronomy) mirror gear drive mechanism; replacing M074 (specimen mass measurement) electronics package; clearing jammed film plate on S183 (ultraviolet panorama); manual opening of stuck S054 (X-ray spectographic telescope) door; camera/magazine changeout for S082 (extreme ultraviolet spectograph/spectroheliograph); clearing particle contamination from S052 (white light coronograph) occulting disc.

The second manned crew installed a larger thermal canopy over the parasol which provided long-duration thermal protection. A complex package of six ancillary rate gyros was installed and provided the needed control and backup control for pointing the Skylab. The second crew pressurized and flushed the condensate lines and replaced the discharge assembly, thus allowing normal use of the waste water management system. Detailed inspection and checkout of the primary and secondary airlock module coolant loops to determine internal and external leakage were performed, and the information relayed to ground for analysis.

The second crew also corrected significant experiment and support system problems, such as jettisoning T027 (Apollo telescope mount contamination measurement) to clear scientific airlock; replacing video tape recorder and teleprinter head; removing ATM door ramps to improve door closing and opening operations; repairing Mark I exerciser; correcting S019 (ultraviolet stellar astronomy) articulating mirror malfunction; checking out and restoring lighting control; replacing defective TV monitor and TV power cable; replacing and analyzing causes for malfunctioning tape recorders; correcting seal problem on M092 (inflight lower body negative pressure); hooking up the sensor down-link data cable assembly of the Earth resources experiment package for ground data diagnosis.

The third manned crew successfully replaced the malfunctioning ATM TV monitor in the control and display panel; this restored the display redundancy critical for effective ATM solar pointing. The depleted primary AM coolant loop was reserviced using the coolanol servicing kit resupplied on SL-4, thus ensuring continued thermal control of a critical internal Skylab system. An automatic timer and cable was launched on SL 4 and installed by the crew on the ATM control and display console to correct erratic exposure operation resulting in loss of S082 experiment data.

During SL-4 EVA 1, the crew performed troubleshooting and inspection of the inoperative S193 (microwave radiometer/scatterometer radar and altimeter) antenna and then locked the antenna in a position to permit continued operation of the experiment. The crew installed replacement units and provided an operational TV system to replace the system which failed on SL-3. The 'noisy' multispectral scanner hardware was corrected by replacing a modified attenuator and the sensitivity of the unit was improved by installing a modified detector/coder/dewar assembly. Successful crew investigation and rerouting of dumpline into the waste tank eliminated venting disturbances.

Other experiment and support system repairs performed by the third crew included installation of liquid crystal thermometers for rate gyro 6-pack assembly; 4 minutes 24 seconds, during which the nine crewmen accrued a total of 12 351 hours 43 minutes 12 seconds of flight time.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use